### v1

@@ 4. Proof of Theorems A and B
This section is devoted to prove our main results. Before proving Theorem A, we briefly outline the main idea behind the argument. The strategy combines the density of periodic orbits, in non-trivial sectional-hyperbolic chain-recurrence classes, and the denseness of its stable and unstable manifolds. In this way, for proving Theorem A we need to prove the following lemma:
@@
Now, since $M$ is a compact manifold, there is a well-defined injectivity radius $e_{0}>0$ for the exponential map $\exp_{x}$, for every $x\in M$. For every regular point $x\in C(\sigma_{Y})$ and $\delta>0$, we say that $D_{\delta}(x)\subset M$ is a center-unstable section disc (or simply cu-section disc) of $Y$, if it is given by $\exp_{x}(B^{c-1}_{\delta}(0))$, where $B^{c-1}_{\delta}(0)\subset T_{x}M$ is the c1c-1-dimensional ball centered at $x$ and radius $\delta$, and $T_{y}D_{\delta}(x)\oplus\langle Y(y)\rangle\subset\mathscr{C}_{y}^{c}(a)$ for every $y\in D_{\delta}(x)$.
Let $e_{0}>0$ be the injectivity radius of $\exp$. Fix $0<\varepsilon<e_{0}$ so that $B_{\varepsilon}(C(\sigma_{Y}))\subset U$, where $Y\in\mathcal{U}$ and $U\subset M$ are given by Theorem 3.1. Given $T>0$, we can shrinking $U$ in a such way that $Y_{[0,T]}(x)\in U$. Hence, in a similar way that of [13, Section 4.2], there are finitely many, and well defined, cross sections $D_{Y}\subset M$, which are uniformly away from the singularities of $C(\sigma_{Y})$ for every $Y\in\mathcal{U}$. The next lemma shows that these cu-section discs expands, in a certain way, along the orbit of points of $C(\sigma_{Y})$.
@@
Let $e_{0}>0$ be the injectivity radius of $\exp$. Fix $0<\varepsilon<e_{0}$ so that $B_{\varepsilon}(C(\sigma_{Y}))\subset U$, and let consider $0<\delta<\varepsilon$. Then, for every regular point $x\in C(\sigma_{Y})$ there is $t_{x}^{\prime}>0$ such that $Y_{t_{x}^{\prime}}(x)$ belongs to a cross-section $S_{x}$. In particular, there are $\alpha_{1},\alpha_{2},\beta>0$ such that $\alpha_{1}\leq|Y(Y_{t_{x}^{\prime}}(x))|\leq\alpha_{2}$, and the angle between $Y(Y_{t_{x}^{\prime}}(x))$ and $S_{x}$ is bigger than $\beta$. Therefore, we have $DY_{t}(\mathscr{C}^{c}(y))\subset\mathscr{C}^{c}(Y_{t}(y))$, for every $y\in D_{\delta}(x)$ and every $t>0$ satisfying $Y_{t}(y)\subset U$. Hence, $D_{r_{t}}(Y_{t}(x))=Y_{t}(D_{\delta}(x))\cap\exp_{Y_{t}(x)}B_{\varepsilon}(0)$, where $t>0$ and $0<r_{t}\leq\varepsilon$, is a $cu$-section disc of $Y$. In other words, the $cu$-section discs are invariant for the flow.
@@
Next, we are ready to prove Theorem A.
Proof of Theorem A.
Consider the residual set $\mathcal{R}$ given by Lemma 3.2. Let $X\in\mathcal{R}$ and let $C(\gamma)$ be a non-trivial sectional-hyperbolic chain-recurrent of $X$.
By Lemma 3.5 we have that $C(\gamma)$ is a homoclinic class. Consequently, the stable manifold of $\gamma$ is dense in $C(\sigma)$ and meets transversally every submanifold of dimension $\dim(F)$ that is tangent to $\mathscr{C}^{c}$, has inner radius $\varepsilon$, and intersects a sufficiently small neighborhood $U^{\prime}\subset U$ of $C(\sigma)$. Then, following the arguments in the proof of [13, Theorem 4.1], we can see that this property holds for vector fields $Y$ in a small $C^{1}$-neighborhood $\mathcal{U}_{X}$ of $X$. By Theorem 3.1 we have that $C(\sigma_{Y})$ contains periodic orbits. Moreover, for every $\gamma^{\prime}\subset C(\sigma_{Y})$ and a neighborhood $U_{\gamma^{\prime}}\subset U^{\prime}$ of $\gamma^{\prime}$, one has that $W_{\varepsilon}^{u}(\gamma^{\prime})$ is a submanifold tangent to $\mathscr{C}^{c}$ with inner radius $\varepsilon$. Thus, by Lemma 4.1 we have $W^{s}(\gamma_{Y})\cap W_{\varepsilon}^{u}(\gamma^{\prime})\neq\emptyset$. In particular, this shows that the set of transverse intersections $W^{s}(\gamma_{Y})\pitchfork W^{u}(\gamma^{\prime})$ is dense in $W^{u}(\gamma^{\prime})$.
Therefore, if $x\in C(\sigma_{Y})$ and $V_{x}\subset U$ is a neighborhood of $x$, by Lemma 4.1 there is a point $y\in W^{u}(\gamma^{\prime})$ for some periodic orbit $\gamma^{\prime}\subset C(\sigma_{Y})$, which in turns is accumulated by elements of $W^{s}(\gamma_{Y})\pitchfork W^{u}(\gamma^{\prime})$. In other words, $C(\gamma)$ is robustly a homoclinic class. This proves the result. ∎
Once we have proved Theorem A, let us provide the reader with the proof of Theorem B.
Proof of Theorem B.
To begin with, notice that the star property is a $C^{1}$-open property, so that $\mathcal{X}_{*}^{1}(M)$ is a $C^{1}$-open subset of $\mathcal{X}^{1}(M)$. Consequently, $\mathcal{R}^{\prime}=\mathcal{R}\cap\mathcal{X}^{1}_{*}(M)$ is a residual subset of $\mathcal{X}^{1}(M)$, where $\mathcal{R}$ is the residual set given by Theorem A.
On the other hand, by [28, Theorem A], there is a residual subset $\mathcal{S}^{\prime}\subset\mathcal{X}^{1}_{*}(M)$ such that for any $Y\in\mathcal{S}^{\prime}$ and $\Lambda$ is a non-trivial chain-recurrent class, one of the following holds:
  1. (1)
    $h_{top}(X|_{\Lambda})>0$, $\Lambda$ contains a periodic orbit and is isolated; or
  2. (2)
    $h_{top}(\Lambda)=0$ and $\Lambda$ is sectional-hyperbolic aperiodic class.
By setting $\mathcal{S}=\mathcal{R}^{\prime}\cap\mathcal{S}^{\prime}$, any non-trivial sectional-hyperbolic chain-recurrent class for $Y\in\mathcal{S}$ contains a periodic orbit by Theorem A. Hence, option 2 cannot occur, and the proof is complete. ∎
Acknowledgments. The authors would like to thank Professor Alexander Arbieto for pointing a gap in the proof of a crucial lemma in an earlier version of this work and for valuable suggestions that helped us to figure out a solution for this gap. K. J. Vivas was supported by ANID Proyecto FONDECYT Iniciación 11250633, Chile.

### v2

@@ 4. Proof of the Main Theorem
This section is devoted to prove our main results. Before proving the Main Theorem, we briefly outline the main idea behind the argument. The strategy combines the density of periodic orbits in non-trivial sectional-hyperbolic chain-recurrence classes, and the denseness of its stable and unstable manifolds. In this way, for proving the Main Theorem we need to prove the following lemma:
@@
For the next lemmas we need to recall another type of cross-section that were defined in [14]. These cross-section are different from the cross-sections introduced in Section 3 and hence here we call them adapted cross-sections to avoid confusion. Recall that since $X$ is a generic vector field, one can assume that $X$ has finitely many hyperbolic singularities. Thus, we can suppose the same follows for the any vector field in $\mathcal{U}$. Following [14, Section 4.2] for any singularity singularity $\sigma\in C(\sigma_{Y})$, there is a co-dimension one submanifold $S_{\sigma}$ so that for every $\sigma\in C(\sigma_{Y})$ it holds:
  • The submanifolds $\overline{S_{\sigma}}$ are pairwise disjoint
  • $\overline{S_{\sigma}}\cap Sing(Y)=\emptyset$ and the angle between $Y(x)$ and $T_{x}S_{\sigma}$ is bigger than $\alpha_{\sigma}>0$, for every $x\in S_{\sigma}$.
By considering $\alpha=\min\{\alpha_{\sigma};\sigma\in Sing(Y)\}$ and $S_{Y}=\cup_{\sigma\in Sing(Y)}S_{\sigma}$ it holds that there is a compact set $\Delta\subset S_{Y}$ such that
  • $\overline{S_{Y}}\cap Sing(Y)=\emptyset$ and the angle between $Y(x)$ and $T_{x}S_{Y}$ is bigger than $\alpha>0$, for every $x\in S_{Y}$.
  • The interior of $\Delta_{Y}$ intersects the any forward orbit in $M\setminus\cup_{\sigma\in Sing(Y)}W^{s}_{loc}(\sigma)$.
  • For every $\sigma\in Sing(Y)$, $S_{Y}\cap W^{s}_{loc}(\sigma)$ and $S_{Y}\cap W^{u}(\sigma)$ are a one-codimensional spheres inside $W^{s}_{loc}(\sigma)$ and $W^{u}_{loc}(\sigma)$, respectively.
Hereafter, the sets $S_{\sigma}$ and $S_{Y}$ will be called the adapted cross-sections of $Y$.
Now, since $M$ is a compact manifold, there is a well-defined injectivity radius $e_{0}>0$ for the exponential map $\exp_{x}$, for every $x\in M$. For every regular point $x\in C(\sigma_{Y})$ and $\delta>0$, we say that $D_{\delta}(x)\subset M$ is a center-unstable section disc (or simply cu-section disc) of $Y$, if it is given by $\exp_{x}(B^{c-1}_{\delta}(0))$, where $B^{c-1}_{\delta}(0)\subset T_{x}M$ is the c1c-1-dimensional ball centered at $x$ and radius $\delta$, and $T_{y}D_{\delta}(x)\oplus\langle Y(y)\rangle\subset\mathscr{C}_{y}^{c}(a)$ for every $y\in D_{\delta}(x)$. Let $e_{0}>0$ be the injectivity radius of $\exp$. Fix $0<\varepsilon<e_{0}$ so that $B_{\varepsilon}(C(\sigma_{Y}))\subset U$, where $Y\in\mathcal{U}$ and $U\subset M$ are given by Theorem 3.1. Given $T>0$, we can shrinking $U$ in a such way that $Y_{[0,T]}(x)\in U$. The next lemma shows that these cu-section discs expands, in a certain way, along the orbit of points of $C(\sigma_{Y})$.
@@
Let $e_{0}>0$ be the injectivity radius of $\exp$. Fix $0<\varepsilon<e_{0}$ so that $B_{\varepsilon}(C(\sigma_{Y}))\subset U$, and let consider $0<\delta<\varepsilon$. Let $S_{Y}$ be an adapted cross-section for $Y$. Thus, for every regular point $x\in C(\sigma_{Y})$ there are $\sigma\in Sing(X)$ $S_{\sigma}$ and $t_{x}^{\prime}>0$ such that $Y_{t_{x}^{\prime}}(x)\in S_{\sigma}=S_{x}$. In particular, there are $\alpha_{1},\alpha_{2},\beta>0$ such that $\alpha_{1}\leq|Y(Y_{t_{x}^{\prime}}(x))|\leq\alpha_{2}$, and the angle between $Y(Y_{t_{x}^{\prime}}(x))$ and $S_{x}$ is bigger than $\beta$. Therefore, we have $DY_{t}(\mathscr{C}^{c}(y))\subset\mathscr{C}^{c}(Y_{t}(y))$, for every $y\in D_{\delta}(x)$ and every $t>0$ satisfying $Y_{t}(y)\subset U$. Hence,
$D_{r_{t}}(Y_{t}(x))=Y_{t}(D_{\delta}(x))\cap\exp_{Y_{t}(x)}B_{\varepsilon}(0),$
where $t>0$ and $0<r_{t}\leq\varepsilon$, is a $cu$-section disc of $Y$. In other words, the $cu$-section discs are invariant for the flow.
@@
Next, we are ready to prove the Main Theorem.
Proof of the Main Theorem.
Consider the residual set $\mathcal{R}$ given by Lemma 3.2. Let $X\in\mathcal{R}$ and let $\Lambda$ be a non-trivial sectional-hyperbolic chain-recurrent of $X$ which is not reduced to either a homoclinic loop or a saddle connection between singularities.
By Theorem 3.1 we have that $\Lambda$ is a homoclinic class. Consequently, there is a periodic orbit $\gamma\subset\Lambda$ whose stable manifold is dense in $\Lambda$ and meets transversally every submanifold of dimension $\dim(F)$ that is tangent to $\mathscr{C}^{c}$, has inner radius $\varepsilon$, and intersects a sufficiently small neighborhood $U^{\prime}\subset U$ of $\Lambda$. Then, following the arguments in the proof of [14, Theorem 4.1], we can see that this property holds for vector fields $Y$ in a small $C^{1}$-neighborhood $\mathcal{U}_{X}$ of $X$. By Theorem 3.1 we also have that $\Lambda_{Y}$ contains periodic orbits. Moreover, for every $\gamma^{\prime}\subset\Lambda_{Y}$ and a neighborhood $U_{\gamma^{\prime}}\subset U^{\prime}$ of $\gamma^{\prime}$, one has that $W_{\varepsilon}^{u}(\gamma^{\prime})$ is a submanifold tangent to $\mathscr{C}^{c}$ with inner radius $\varepsilon$. Thus, by Lemma 4.1 we have $W^{s}(\gamma_{Y})\cap W_{\varepsilon}^{u}(\gamma^{\prime})\neq\emptyset$. In particular, this shows that the set of transverse intersections $W^{s}(\gamma_{Y})\pitchfork W^{u}(\gamma^{\prime})$ is dense in $W^{u}(\gamma^{\prime})$.
Therefore, if $x\in\Lambda_{Y}$ and $V_{x}\subset U$ is a neighborhood of $x$, by Lemma 4.1 there is a point $y\in W^{u}(\gamma^{\prime})$ for some periodic orbit $\gamma^{\prime}\subset\Lambda_{Y}$, which in turns is accumulated by elements of $W^{s}(\gamma_{Y})\pitchfork W^{u}(\gamma^{\prime})$. In other words, $\Lambda_{Y}$ is robustly a homoclinic class. This proves the result. ∎
Acknowledgments. The authors would like to thank some colleagues for valuable comments that helped them to improve the presentation of this work. Precisely, we thank Professor Alexander Arbieto and Professor Fan Yang for the careful reading of this manuscript and for pointing out several issues in an earlier version of this work. K. J. Vivas was supported by ANID Proyecto FONDECYT Iniciación 11250633, Chile.